OpenStudy (anonymous):

If i = sq. root of -1 and a and b are non-zero real numbers, what is 1/ (a+bi) ?

OpenStudy (anonymous):

Multiply by conjugate.

OpenStudy (anonymous):

(a-bi)/(a-bi) right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Multiply by that.

OpenStudy (anonymous):

I got something weird like this (a-bi)/a^2-b^2i^2

OpenStudy (anonymous):

In the answer choices, theres no i

terenzreignz (terenzreignz):

That's not weird, just... cumbersome \[\huge \frac{a-bi}{a^2-b^2i^2}\]But what is \(\Large i^2 = ?\)

OpenStudy (anonymous):

oh i see... i^2=-1

terenzreignz (terenzreignz):

So... it becomes...?

OpenStudy (anonymous):

It is really really really really important to remember \(i^2=-1\) and \(i=\sqrt{-1}\) It's the whole reason we multiply by conjugate.

OpenStudy (anonymous):

(a-bi) / a^2+b^2

terenzreignz (terenzreignz):

And you're done.

OpenStudy (anonymous):

\[ \left(\frac{a}{a^2+b^2}\right) - \left(\frac{b}{a^2+b^2}\right)i \]

OpenStudy (anonymous):

huh?

terenzreignz (terenzreignz):

He just separated the real part from the imaginary.

OpenStudy (anonymous):

ok